### v2

@@ 5.3 KCQ Hamiltonian and Eigenstates
The difference between the coherent state $|-\alpha\rangle$ and the ground state $|\phi\rangle$ of the detuned-KCQ when $\Delta=2K$ is shown in Fig. 10. The Wigner functions of the ground states when $\epsilon_{2}/K=1$ and $\epsilon_{2}/K=4$ are also plotted in the inset indicated by the orange and blue stars. Both the results from perturbation theory and numerical diagonalization indicate the rapidly decreasing difference between the coherent state and the detuned-KCQ ground state. At the operating point in this work, such a difference is below 0.3%.
5.4 Device Fabrication
@@
The device in this work is based on superconducting circuits patterned on an intrinsic silicon (Si) wafer with a resistivity $\rho\geq 10\mathrm{k\Omega\cdot cm}$. The fabrication process is briefly depicted in the flow chart shown in Fig. 11a. Before the metal deposition, the wafer is cleaned by Piranha solutions ($\mathrm{H_{2}SO_{4}+H_{2}O_{2}}$) and hydrofluoric acid (HF) to remove the organic contamination and silicon oxide on the surface. The wafer is then transferred to a vacuum box to a Kert J. Lesker sputtering tool for niobium (Nb) deposition. By adjusting the sputtering temperature and pressure, a uniform niobium layer with compressive stress and high superconducting transition temperature[62] is sputtered onto the wafer. The superconducting transition temperature of the film is measured to be 8.94 K, as shown in Fig. 11b. The large structures, including the microwave buses, filter, quarter-wavelength resonators and capacitor pads, are patterned on the Nb film using photolithography and reactive ion etching (RIE). After a second cleaning process with HF to further remove oxides on Si and Nb surfaces, the Josephson junctions are defined with e-beam lithography (EBL) and bridge-free two-angle shadow evaporation of aluminum (Al) films[63]. Electrical contact between the Al Josephson junctions and the Nb capacitor pads is formed by the bandaid process with argon ion-milling[64]. Once diced, the device is packaged into a oxygen-free copper box for testing in the dilution refrigerator.
@@
5.5 GST Circuit Design
GST is a self-consistent calibration free protocol for characterizing the noisy implementation of a informationally complete set of quantum gates[65, 66, 67]. GST can estimate the Pauli transfer matrix (PTM) of quantum gates in a SPAM-free manner, and is capable of fitting models with physical constraints such as completely positive, trace preserving (CPTP) constraints[65, 66, 67]. This technique has been applied to various platforms of quantum computing, including ion traps[68], spin qubits[69, 70] and superconducting circuits[71]. GST usually executes a set of circuits that consist of a state preparation and readout sequence (fiducials) and repeated sequences of gate operations (germs) that amplify specific types of error [72, 65].
Because the local gates of KCQ are the discrete $X(\pi/2)$ gate and continuous $Z(\theta)$ gates, we construct the informationally complete set of quantum gates as $\{X(\pi/2),Z(\pi/2),Y(\pi/2)\}$, where the $Y(\pi/2)$ gate is complied using local gates as $X(\pi)Z(\pi/2)X(\pi/2)Z(\pi/2)$. Each GST circuit has the form $F_{i}G^{l}F_{j}$, where $F_{i}$ and $F_{j}$ are fiducial sequences, $G$ is a germ sequence, and $l$ is the maximum integer such that the length of $G^{l}$ is at most some maximum length $L$. Using the pyGSTi python package[73], we generate the GST circuits which have a germ set with $10$ germs of maximum length $5$. The fiducials and germs are shown in Tab. 2.

### v3

@@ Discussion
The difference between the coherent state $|-\alpha\rangle$ and the ground state $|\phi\rangle$ of the detuned-KCQ when $\Delta=2K$ is shown in Fig. 5b. The Wigner functions of the ground states when $\epsilon_{2}/K=1$ and $\epsilon_{2}/K=4$ are also plotted in the inset indicated by the orange and blue stars. Both results from perturbation theory and numerical diagonalization indicate the rapidly decreasing difference between the coherent state and the detuned-KCQ ground state. At the operating point in this work, such a difference is below 0.3%.
Detuned-KCQ Bit-flip Time $T_{z}$ Simulations
We illustrate the simulation of bit-flip times $T_{z}$ in this section and emphasize the enhancement of bit-flip times when the two-photon stabilization drive is detuned by $2K,4K$... due to the degenerate higher excited states (39, 38, 23).
Figure 6: Master equation simulation. a, The bit-flip times $T_{z}$ with different mean photon number and their master equation simulation with the fitting parameters taking different incoherent processes into account. b, detuned-KCQ Hamiltonian spectrum with detuned two-photon stabilization drive. c, The bit-flip times $T_{z}$ with different detuning and their master equation simulation.
The Kerr-cat qubit inevitably disperses irretrievable information to the environment, leading to incoherent processes. We model the environment by a chain of harmonic oscillators with infinite modes and temperature $T_{\rm E}$. While the exact dynamics of arbitrary system-environment interaction and environment evolution are intractable analytically, one can construct the master equation describing various effects of the environment on the system under assumptions that generally hold in the experiments, including weak system-environment coupling and a Markovian thermal environment. For simplicity, we consider an environment with fixed thermal distribution and energy-independent coupling with the system. With the above approximations, the master equation has the following form,
where $\hat{\cal O}_{l}$ describe three sets of incoherent processes, including $\hat{\cal O}_{1,\downarrow}\leavevmode\nobreak\ (\hat{\cal O}_{1,\uparrow})=\hat{a}\leavevmode\nobreak\ (\hat{a}^{\dagger})$ for the single-photon dissipation (excitation) process, $\hat{\cal O}_{2,\downarrow}\leavevmode\nobreak\ (\hat{\cal O}_{2,\uparrow})=\hat{a}^{2}\leavevmode\nobreak\ (\hat{a}^{\dagger^{2}})$ for the two-photon dissipation (excitation) process, and $\hat{\cal O}_{\phi}=\hat{a}^{\dagger}\hat{a}$ for the dephasing process. The corresponding coefficients $\gamma_{1,\downarrow}(\gamma_{1,\uparrow})\propto\kappa^{(1)}$, $\gamma_{2,\downarrow}(\gamma_{2,\uparrow})\propto\kappa^{(2)}$, and $\gamma_{\phi}\propto\kappa_{\phi}$ are single-photon dissipation (excitation), two-photon dissipation (excitation), and dephasing rates, where $\kappa^{(1)}$ and $\kappa^{(2)}$ are the single-photon and two-photon spontaneous decay rates, respectively. A detailed derivation of the master equation along with the expressions for $\gamma_{1,\downarrow}(\gamma_{1,\uparrow}),\gamma_{2,\downarrow}(\gamma_{2,\uparrow})$ and $\gamma_{\phi}$ is presented in references (44, 23, 33). Beyond the three incoherent processes above, we also consider a stochastic process that leads to fluctuations in the SNAIL nonlinear oscillator frequency with a noise strength of $\xi$ (23, 33).
Figure 7: Detuned-KCQ readout, initialization and gate calibration. a, The histogram of CQR. b, The collapse of wave-functions corresponding to CQR. c, The QNDness optimization of CQR with various readout strengths. d, The SNAIL nonlinear oscillator spectrum in the rotating frame as a function of detuning. e, The sequence for calibrating $X(\pi/2)$ gate with 400-$\mathbf{\mu}s$ relaxation time in the initialization. f, The tunneling between the ground states when the relaxation time is long enough in the initialization. g, The measured $|\pm Z\rangle$ population oscillation. h, The numerical simulation of the results above it. i, The sequence for observing the excited states tunneling with 4-$\mathbf{\mu}s$ relaxation time in initialization. j, The tunneling between the excited states when the relaxation time is short. k The measured population oscillation between excited states. l, The numerical simulation of the results above.
With such four processes in a thermal environment with thermal photon population $n_{\text{th}}=4\%$, and parameters $\kappa^{(2)}=7\text{MHz}$, $\kappa_{\phi}=100\text{Hz}$, $\xi=40\text{kHz}$, the experimental behavior of the bit-flip times $T_{z}$ is accurately modeled as shown by the agreement between the experiment data and master equation simulation results in Fig. 2. The thermal population corresponds to a temperature of 86.9 mK, which is higher than the base temperature of the dilution fridge due to the strong microwave drives. The pure dephasing rate $\kappa_{\phi}$ is lower than that estimated from the $T_{2}$ values, which is also observed in all related literature (27, 38, 33), and might be due to the non-uniform power spectrum of the dephasing process under strong drive.
Here, we illustrate the effects of each process in detail. As shown in Fig. 6a, we fit the bit-flip times of a resonant-KCQ as a function of two-photon stabilization drive strength $\epsilon_{2}$. When only considering the single-photon dissipation and excitation, i.e. $\kappa^{(2)}=\kappa_{\phi}=\xi=0$, the master equation simulation follows the blue curve, overestimating the bit-flip times $T_{z}$. By introducing the stochastic frequency fluctuation of the SNAIL nonlinear, i.e., $\kappa^{(2)}=\kappa_{\phi}=0$, the master equation simulation successfully modeled the bit-flip times with small $\epsilon_{2}$ following the yellow curve, but failed to model the strongly driven regime. Finally, only by including all four processes does the master equation simulation accurately model the experimental data, as shown by the red curve.
Beyond the influence on bit-flip times $T_{z}$ by the two-photon stabilization drive strength, the detuning of two-photon stabilization drives, with $\Delta=2K,4K$..., can lead to degenerate higher excited states marked by the stars in Fig. 6b, and increase the bit-flip times. We observe $N_{\text{deg}}$ degenerate points, corresponding to $N_{\text{deg}}-1$ higher excited degenerate states, when $\Delta=2(N_{\text{deg}}-1)K$. With a fixed two-photon stabilization drive strength $\epsilon_{2}=4.2K$, we measured the bit-flip times with various detuning $\Delta$, as shown in Fig. 6c. The increase of the bit-flip times when $\Delta=2K$ and $4K$ is observed both in experimental data and predicted in the master equation simulations, demonstrating the validity of the theoretical model.
Quantum control Toolbox for a Detuned-KCQ
In order to operate a system as a qubit for quantum computing, it is crucial to have reliable state readout, universal quantum control and state initialization.
The readout of the detuned-KCQ is enabled by a beam-splitter interaction between the detuned-KCQ mode $\hat{a}$ and the readout resonator mode $\hat{b}$. Due to the detuning of the stabilization drive, the rotating frame frequency is referenced at half of the stabilization drive frequency instead of the qubit frequency. Therefore, by applying a cat-qubit-readout (CQR) drive at frequency $\omega_{CQR}=\omega_{R}-\omega_{S}/2$, we can engineer a beam-splitter interaction given as
Such an interaction will populate the readout resonator mode $\hat{b}$ depending on the expectation value of the detuned-KCQ mode $\hat{a}$ as
To calibrate the CQR drive, we apply two consecutive CQR pulses with a duration of 4.5 $\mathbf{\mu}s$ for each, and define the QNDness as $(P\left[M_{|+\alpha\rangle}|M_{|+\alpha\rangle}\right]+P\left[M_{|-\alpha\rangle}|M_{|-\alpha\rangle}\right])/2$, where $P\left[M_{A}|M_{B}\right]$ represents the probability of getting result $A$ with the second CQR pulse conditioned on having obtained result $B$ with the first CQR pulse. The QNDness is optimized with respect to the CQR drive strengths shown in Fig. 7c, resulting in good separations shown by the histogram in Fig. 7a. Such a quantum non-demolition readout can be utilized to initialize the qubit through heralding. However, as shown in Fig. 7b and discussed later in this section, the qubit states have to stay in the ground states before heralding because such readout cannot distinguish the excited states and the ground state in the same potential well.
The universal single-qubit control of the detuned-KCQ consists of a discrete $X(\pi/2)$ gate and continuous $Z(\theta)$ rotation gates, both implemented via deformation of the double-well pseudo-potential. The $Z(\theta)$ gates are realized using a single-photon drive at frequency $\omega_{S}/2$ to engineer a Hamiltonian $\hat{H}_{d}=\hat{a}^{\dagger}\Omega/2+\hat{a}\Omega^{*}/2$. This Hamiltonian induces an energy splitting of $2\text{Re}[\Omega]\alpha$ between the $|\pm Z\rangle$ states, resulting in Rabi oscillations on the equator of the Bloch sphere. Calibration and numerical simulations of the $Z(\theta)$ gates are detailed in Suppl. Info. Section 3. The $X(\pi/2)$ gate, which corresponds to coherent tunneling between the two pseudo-potential wells, is implemented by adiabatically lowering the energy barrier separating them. We implement such gate by phase-modulating the two-photon stabilization drive $\epsilon_{2}(t)=\epsilon_{2}(0)e^{-ig(t)}$ with $g(t)$ defined as
where $T_{g}$ is the gate time, $\delta_{0}$ is the modulation depth, and $f(t)=\text{exp}\left(-\frac{8(t-Tg/3)^{2}}{T_{g}^{2}}\right)$ follows a Gaussian profile. The asymmetric modulation profile is chosen to minimize the gate error. Such phase modulation adiabatically introduces a blue-detuning to the two-photon stabilization drive and correspondingly, lowers the energy barrier of the pseudo-potential to allow the inter-well tunneling, as shown in Fig. 7f.
We calibrate the $X(\pi/2)$ gate with the sequence shown in Fig. 7e. A CQR pulse (labeled “Herald”) initializes the qubit into $|+Z\rangle$ state through heralding. Then, two $X(\pi/2)$ gate pulses with various modulation depth $\delta_{0}$ and duration are applied. Finally, another CQR pulse (labeled “Readout”) is applied to read out the qubit state. The measured expectation of the Pauli Z operator $\langle\hat{Z}\rangle$ is shown in Fig. 7g, which clearly shows population oscillating coherently between the $|\pm Z\rangle$ states. The numerical simulation, based on solving the time-dependent Schrödinger equations, is shown in Fig. 7h in good agreement with the experimental results. Therefore, we calibrate the $X(\pi/2)$ gate for a detuned-KCQ by optimizing the gate time and the modulation depth to achieve the largest $\langle\hat{Z}\rangle$ conversion. Importantly, a sufficiently long relaxation time (400 $\mathrm{\mu s}$) is necessary for the initialization as discussed in the following.
The initialization of the detuned-KCQ is implemented through the CQR pulse by heralding. The SNAIL nonlinear oscillator mode predominantly remains in the vacuum state $|0\rangle$ with negligible thermal noise at low temperature. Therefore, the resonant-KCQ can be initialized by adiabatically ramping up the two-photon stabilization drive followed immediately by a CQR drive, which maps the vacuum state to the cat state and then collapses it into $|\pm Z\rangle$. However, this technique fails for the detuned-KCQ because the finite detuning changes the energies of SNAIL nonlinear oscillator states in the rotating frame. As shown in Fig. 7d, a finite $\Delta$ alters the spectrum in the rotating frame, causing $|0\rangle$ to no longer be the ground state, while the cat states remain the ground states of the detuned-KCQ Hamiltonian. Consequently, when the two-photon stabilization drive is ramped up, the $|0\rangle$ state is mapped to the higher excited states instead of the ground states. The subsequent CQR drive will then collapse the qubit into excited states in one of the double wells of the pseudo-potential, resulting in a failure to initialize the qubit into $|\pm Z\rangle$ states. To address this, it is crucial to introduce sufficient relaxation time before applying the CQR drive, allowing the system to decay into the degenerate ground states. This ensures the validity of preparing the $|\pm Z\rangle=|\pm\alpha\rangle$ state through heralding.
To observe the effect of higher excited states, we applied a sequence similar to that used for calibrating the $X(\pi/2)$ gate to a detuned-KCQ system with $\Delta=2K$, but with the relaxation time reduced to $5\leavevmode\nobreak\ \mathcal{\mu}s$, as shown in Fig. 7i. Due to this shortened relaxation time, the detuned-KCQ states predominantly occupy the excited states, and the pulse sequence captures the population oscillation between these excited states within each well, as illustrated in Fig. 7j. The experimental results, presented in Fig. 7k, show a reduced contrast and tunneling occurring at a lower modulation depth due to the low lifetime and high energy of the excited states, which is in excellent agreement with the numerical simulation shown in Fig. 7l.
[…]
B.Q., Ah.H. and K.W. led and organized the project. B.Q. fabricated the device, performed the measurement and analyzed the data. Ah.H. and G.K. designed the device. Ah.H, G.K., and L.B.N. set up the cryogenic apparatus and initialized the measurement. I.H. and B.B. performed the theoretical analysis. Ah.H., K.W., N.J., N.G. and N.F. assisted with the measurement. J.H. offered the GST circuit and analyzed the GST data. Ak.H. assisted with the GST data analysis. Z.P., L.C., Z.K., C.J., H.K., and K.L. assisted with the device fabrication. J.D., A.J., D.S. and I.S. supervised the project and offered necessary supports. B.Q., K.W. and L.B.N. wrote the manuscript with input from all authors.
[…]
Supplementary Information
1 Measurement Hardware, Device Layout and Device Parameters
The Kerr-cat qubit is implemented on a planar superconducting circuit device. The equivalent circuit diagram of the device with the measurement wiring diagram is shown in Fig. 8a, and the false-color micrograph of the device chip is shown in Fig. 8b. The key component of the Kerr-cat qubit device, SNAIL nonlinear oscillator (yellow), consists of two SNAILs shunted by a coplanar capacitor, where each SNAIL is formed by three large Josephson junctions connected in parallel with a small Josephson junction, as shown in Fig, 8c. A readout resonator (dark green) is capacitively coupled to the SNAIL nonlinear oscillator with a dispersive shift $\chi$ to enable the quantum nondemolition readout of the Kerr-cat qubit, and a Purcell filter (light green) is introduced between the readout resonator and the readout port (pink) to engineer the desired readout resonator decay rate $\kappa_{r}$ with suppressed Purcell decay. We engineered two microwave ports to deliver the microwave efficiently without introducing too much Purcell decay. On the one hand, a control port (orange) is weakly coupled to the SNAIL nonlinear oscillator to deliver the single-photon drive for the continuous $Z(\theta)$ gates. On the other hand, a pump port (purple) with a bandblock filter is strongly coupled to the SNAIL nonlinear oscillator to deliver the two-photon stabilization drive and the cat-qubit readout (CQR) drive, where the bandblock filter strongly suppressed the Purcell decay from this port by blocking the signal around the qubit frequency. Detailed description of the bandblock filter can be found in our previous work(23).
We characterize the Kerr-cat qubit with the microwave network shown in Fig. 8a, where we neglect some filters, amplifiers and attenuators for simplicity and clarity, and a detailed wiring diagram can be found in our previous work(23). The modulation and demodulation of the microwave drive is implemented by two local oscillators, frequency mixers, and a frequency doubler. A readout local oscillator (LO) at readout resonator frequency $\omega_{R}$ is introduced to demodulate the readout signal from the readout port (pink) of the device, and a qubit LO at SNAIL nonlinear oscillator frequency $\omega_{Q}$ is split into three parts to generate the control, readout and stabilization drives. The first one is directly fed into the IQ mixer to generate the single-photon drive, which is connected to the weakly coupled port (orange) of the device. The second one is fed into a frequency doubler to make its frequency doubled to $\omega_{S}\approx 2\omega_{Q}$ and generate the two-photon stabilization drive with the IQ mixer. Finally, the last one is mixed with the readout LO to generate the signal at the cat-qubit readout frequency $\omega_{CQR}\approx\omega_{R}-\omega_{S}/2$. The two-photon stabilization drive and the cat-qubit readout drive are both connected to the strongly coupled port (purple) of the device through a directional coupler.
Table 1: Device Parameters and Calibration Methods
Parameter Value Methods
SNAIL oscillator Kerr $K$ $2\pi\times 1.2$ MHz two-tone spectroscopy
SNAIL large junction $E_{J}$ $2\pi\times 263.2$ GHz room-temp. resistance msmt.(62)
SNAIL small-to-large junction ratio $\beta$ 0.1 room-temp. resistance msmt.(62)
SNAIL oscillator dephasing time $T_{2}$ 3.2 $\mathrm{\mu}s$ time-domain Ramsey msmt.(63)
SNAIL oscillator thermal population $n_{\text{th}}$ 5% $T_{z}$ fitting with theoretical model
2 KCQ Hamiltonian Parameter Calibration
We first extract the KCQ nonlinear oscillator parameters without the two-photon stabilization drive, i.e. $\epsilon_{2}=0$. The KCQ nonlinear oscillator frequency $\omega_{Q}=\omega_{01}$ is the transition frequency between $|0\rangle$ and $|1\rangle$, determined using standard two-tone spectroscopy(64). The Kerr nonlinearity is then extracted by preparing the nonlinear oscillator in the first excited states $|1\rangle$ followed by a similar two-tone spectroscopy. where we observed two peaks corresponding to the $|1\rangle\rightarrow|0\rangle$ transition with frequency $\omega_{01}$ and $|1\rangle\rightarrow|2\rangle$ transition with frequency $\omega_{12}=\omega_{01}-2K$, which allows us to extract the Kerr nonlinearity $K=1.2\leavevmode\nobreak\ \text{MHz}$, as shown in Fig. 9a.
Subsequently, we apply the two-photon stabilization drive to create cat states in the SNAIL nonlinear oscillator, and extract the value of $\epsilon_{2}$. With the two-photon stabilization drive, the KCQ Hamiltonian supports cat states $|C_{\alpha}^{\pm}\rangle\propto|+\alpha\rangle\pm|-\alpha\rangle$ with mean photon number $|\alpha|^{2}=(\epsilon_{2}+\Delta/2)/K$. Because $K$ and $\Delta$ are already calibrated, we only need to calibrate $|\alpha|^{2}$. As mentioned in reference(24, 23), a single photon drive $H_{d}=\frac{\Omega}{2}\hat{a}^{\dagger}+\frac{\Omega^{*}}{2}\hat{a}$ induces a transition between the two cat states with the rate $\Omega_{c}=2\text{Re}[\Omega]\alpha$ with $\epsilon_{2}>0$. Conversely, without the two-photon stabilization drive ($\epsilon_{2}=0$), the same drive can induce the transition between $|0\rangle$ and $|1\rangle$ states of the nonlinear oscillator with the rate $\Omega$. Therefore, by tuning the phase of the single photon drive to make $\epsilon_{1}$ real and measuring the two transition rates, we can extract, $\alpha=\Omega_{c}/(2\text{Re}[\Omega])$, as shown in Fig. 9b.
3 Detuned-KCQ Continuous $Z(\theta)$ Gates Calibration
The continuous $Z(\theta)$ gates is engineered by a single-photon drive at half of the stabilization drive frequency, which shifts the energy of the two pseudo-potential well, as shown in Fig. 10a. We calibrate $Z(\theta)$ gates following the pulse sequence shown in Fig. 10b. The pre-calibrated $X(\pi/2)$ gates are used to prepare and measure the state along Y axis. The phase and duration of the single-photon drive are swept to obtain the Chevron-like plot shown in Fig. 10d, which is in good agreement with the numerical simulation shown in Fig. 10c. The $Z(\theta)$ gate speed can be increased by increasing either the cat qubit size or the single-photon drive strength. However, the large single-photon drive strength will lead to non-adiabatic error discussed in the Main Text.
Figure 10: Calibration of $Z(\theta)$ gates. a The shifted energy profile when applying single-photon drive. b The sequence for calibration the $Z(\theta)$ gates. c Simulation of the measured Chevron-like plot. d The measured Chevron-like plot with the sequence above.
4 Device Fabrication
@@
The device in this work is based on superconducting circuits patterned on an intrinsic silicon (Si) wafer with a resistivity $\rho\geq 10\mathrm{k\Omega\cdot cm}$. The fabrication process is briefly depicted in the flow chart shown in Fig. 11a. Before the metal deposition, the wafer is cleaned by Piranha solutions ($\mathrm{H_{2}SO_{4}+H_{2}O_{2}}$) and hydrofluoric acid (HF) to remove the organic contamination and silicon oxide on the surface. The wafer is then transferred to a vacuum box to a Kert J. Lesker sputtering tool for niobium (Nb) deposition. By adjusting the sputtering temperature and pressure, a uniform niobium layer with compressive stress and high superconducting transition temperature(65) is sputtered onto the wafer. The superconducting transition temperature of the film is measured to be 8.94 K, as shown in Fig. 11b. The large structures, including the microwave buses, filter, quarter-wavelength resonators and capacitor pads, are patterned on the Nb film using photolithography and reactive ion etching (RIE). After a second cleaning process with HF to further remove oxides on Si and Nb surfaces, the Josephson junctions are defined with e-beam lithography (EBL) and bridge-free two-angle shadow evaporation of aluminum (Al) films(66). Electrical contact between the Al Josephson junctions and the Nb capacitor pads is formed by the bandaid process with argon ion-milling(67). Once diced, the device is packaged into a oxygen-free copper box for testing in the dilution refrigerator.
@@
5 GST Circuit Design
GST is a self-consistent calibration free protocol for characterizing the noisy implementation of a informationally complete set of quantum gates(46, 68, 69). GST can estimate the Pauli transfer matrix (PTM) of quantum gates in a SPAM-free manner, and is capable of fitting models with physical constraints such as completely positive, trace preserving (CPTP) constraints(46, 68, 69). This technique has been applied to various platforms of quantum computing, including ion traps(70), spin qubits(71, 72) and superconducting circuits(73). GST usually executes a set of circuits that consist of a state preparation and readout sequence (fiducials) and repeated sequences of gate operations (germs) that amplify specific types of error (74, 46).
Because the local gates of KCQ are the discrete $X(\pi/2)$ gate and continuous $Z(\theta)$ gates, we construct the informationally complete set of quantum gates as $\{X(\pi/2),Z(\pi/2),Y(\pi/2)\}$, where the $Y(\pi/2)$ gate is complied using local gates as $X(\pi)Z(\pi/2)X(\pi/2)Z(\pi/2)$. Each GST circuit has the form $F_{i}G^{l}F_{j}$, where $F_{i}$ and $F_{j}$ are fiducial sequences, $G$ is a germ sequence, and $l$ is the maximum integer such that the length of $G^{l}$ is at most some maximum length $L$. Using the pyGSTi python package(75), we generate the GST circuits which have a germ set with $10$ germs of maximum length $5$. The fiducials and germs are shown in Tab. 2.